sue an experienced shipping clerk can fill a certain order in four hours. tome a new clerk needs seven hours to do the same job. working together how long will it take them to fill the order?
Answer by Cromlix(4381)

1/4 + 1/7 = 1/x rate
Common denominator = 28x
7x + 4x = 28
11x = 28
x = 28/11
x = 2.54 hours
x = 2hrs 32mins 44secs.
Hope this helps.
:-)